# Can LLMs perform structured graph reasoning tasks?

### v2

@@ Abstract
Pretrained Large Language Models (LLMs) have demonstrated various reasoning capabilities through language-based prompts alone, particularly in unstructured task settings (tasks purely based on language semantics). However, LLMs often struggle with structured tasks, because of the inherent incompatibility of input representation. Reducing structured tasks to uni-dimensional language semantics often renders the problem trivial. Keeping the trade-off between LLM compatibility and structure complexity in mind, we design various graph reasoning tasks as a proxy to semi-structured tasks in this paper, in order to test the ability to navigate through representations beyond plain text in various LLMs. Particularly, we design 10 distinct problems of graph traversal, each representing increasing levels of complexity, and benchmark 5 different instruct-finetuned LLMs (GPT-4, GPT-3.5, Claude-2, Llama-2 and Palm-2) on the aforementioned tasks. Further, we analyse the performance of models across various settings such as varying sizes of graphs as well as different forms of k-shot prompting. We highlight various limitations, biases and properties of LLMs through this benchmarking process, such as an inverse relation to the average degrees of freedom of traversal per node in graphs, the overall negative impact of k-shot prompting on graph reasoning tasks, and a positive response bias which prevents LLMs from identifying the absence of a valid solution. Finally, we introduce a new prompting technique specially designed for graph traversal tasks (PathCompare), which demonstrates a notable increase in the performance of LLMs in comparison to standard prompting techniques such as Chain-of-Thought (CoT).  [^1] The code for reproducing the results of this paper can be found at https://github.com/PalaashAgrawal/LLMGraphTraversal
@@ 1 Introduction
Large language Models (LLMs) exhibit interesting reasoning abilities in various types of reasoning tasks [23], including arithmetic [27], logical [26], semantic [14], and symbolic [25]. Building on top of just raw reasoning abilities, various techniques, such as chain-of-thought (CoT) and self-consistency[22], have been proposed to improve the reasoning abilities of LLMs by enhancing the implicit knowledge in the output. However, in relative terms, graph reasoning is not well explored in the domain of LLMs, even though many important subproblems arise in graph reasoning, including subgraph matching [5] and shortest path connectivity [6]. One of the major challenges of grounding graphs through language is rooted in the fact that graphs inherently cover non-linear connectivity between multi-dimensional elements in a condensed format, and hence, reasoning over graphs requires simultaneous tracking of the state of multiple nodes. Broadly in the context of LLMs, graph reasoning is a proxy for understanding how such models establish relations between different entities in a structured setting, and is worth studying, given the impressive emergent abilities observed in recent LLMs.
@@
The major challenge for LLMs in processing graphs directly arises from the uni-dimensionality of input representations, restricting the scope of reasoning that can be achieved using LLMs [1]. Particularly, LLMs are trained on semantic text as next-token predictors, and therefore cannot process structured data representations  [^2] In this paper, we refer to structured representations as data organized according to a specific schema or format that clearly defines the type, relationship, and arrangement of data elements. This structured format allows for efficient processing using standardized algorithms and tools. In this context, graphs are categorized as structured representations by virtue of node connections. On the other hand, LLMs are designed to process unstructured text representations, in which there is a uni-dimensional semantic relationship. By default, LLMs are not equipped to process structured representations through relationship rules. such as graphs and trees directly, without having to break down the representation into verbally semantic prompts [2]. This process often constrains the selection of tasks, and/or leads to loss of complexity. Hence, efficiently capturing the true underlying structure of graphs is particularly challenging for LLMs.
@@
In this paper, we are particularly interested in studying fundamental graph understanding and reasoning abilities of LLMs. While various previous works have studied graph reasoning using LLMs by means of simplified graph structures as enumerated node adjacency lists and connections [2][3][4], graph tasks such as traversal are inherently reduced to trivial search retrieval problems due to flattened representations. Hence, the ability of LLMs to navigate over more complex and structured representations is still not well explored. We explore this particular question in-depth in this paper, by going one step ahead of previous works. Specifically, we prompt LLMs to reason over semi-structured graph representations, instead of plainly verbalized unstructured representations like previous works, while still maintaining the language compatibility of LLMs.  [^3] In this paper’s context, semi-structured representations involve breaking down a structured representation into uni-directional text, while still maintaining some form of element relationship. In our case, adjacency matrix representations involve a tabular relation set, which requires tracking both the column and row origin of a node. This structure introduces complexity to the reasoning task. Particularly, we are interested in exploring the ability in LLMs to handle such representational complexity.

### v3

@@ Abstract
Pretrained Large Language Models (LLMs) have demonstrated various reasoning capabilities through language-based prompts alone, particularly in unstructured task settings (i.e. tasks purely based on language semantics). However, LLMs often struggle with structured tasks, because of the inherent incompatibility of input representation. Reducing structured tasks to uni-dimensional language semantics often renders the problem trivial. Keeping the trade-off between LLM compatibility and structure complexity in mind, we design various graph reasoning tasks as a proxy to semi-structured tasks in this paper, in order to test LLMs’ ability to use representations beyond plain text. In particularly, we design 10 distinct problems of graph traversal, each representing increasing levels of complexity, and benchmark 5 different instruct-finetuned LLMs (GPT-4, GPT-3.5, Claude-2, Llama-2 and Palm-2) on the aforementioned tasks. Further, we analyse the performance of models across various settings such as varying sizes of graphs, as well as different forms of k-shot prompting. We highlight various limitations, biases and properties of LLMs through this benchmarking process, such as an inverse relation to the average degrees of freedom of traversal per node in graphs, the overall negative impact of k-shot prompting on graph reasoning tasks, and a positive response bias which prevents LLMs from identifying the absence of a valid solution. Finally, we introduce a new prompting technique specially designed for graph traversal tasks (PathCompare), which demonstrates a notable increase in the performance of LLMs in comparison to standard prompting techniques such as Chain-of-Thought (CoT).  [^1] The code for reproducing the results of this paper can be found at https://github.com/PalaashAgrawal/LLMGraphTraversal
@@ 1 Introduction
Large Language Models (LLMs) exhibit interesting reasoning abilities in various types of reasoning tasks [23], including arithmetic [27], logical [26], semantic [14], and symbolic [25]. Building on top of just raw reasoning abilities, various techniques, such as chain-of-thought (CoT) and self-consistency [22], have been proposed to improve the reasoning abilities of LLMs by enhancing the implicit knowledge in the output. However, in relative terms, graph reasoning is not well explored in the domain of LLMs, even though many important subproblems arise in graph reasoning, including subgraph matching [5] and shortest path connectivity [6]. One of the major challenges of grounding graphs through language is rooted in the fact that graphs inherently cover non-linear connectivity between multi-dimensional elements in a condensed format, and hence, reasoning over graphs requires simultaneous tracking of the state of multiple nodes. Broadly in the context of LLMs, graph reasoning is a proxy for understanding relations are established between different entities in a structured setting, and is worth studying, given the impressive emergent abilities observed in recent LLMs.
@@
The major challenge for LLMs in processing graphs directly arises from the uni-dimensionality of input representations, restricting the scope of reasoning that can be achieved using LLMs [1]. In particular, LLMs are trained on semantic text as next-token predictors, and hence cannot process structured representations  [^2] In this paper, we refer to structured representations as data organized according to a specific schema or format that clearly defines the type, relationship, and arrangement of data elements. This structured format allows for efficient processing using standardized algorithms and tools. In this context, graphs are categorized as structured representations by virtue of node connections. On the other hand, LLMs are designed to process unstructured text representations with an inherent uni-dimensional semantic relationship. By default, LLMs are not equipped to process structured representations through relationship rules. such as graphs and trees directly, without having to break down the representation into verbally semantic prompts [2]. This process often constrains the selection of tasks, and/or leads to loss of complexity. Hence, efficiently capturing the underlying structure of graphs is particularly challenging for LLMs.
@@
In this paper, we are particularly interested in studying fundamental graph understanding and reasoning abilities of LLMs. While various previous works have studied graph reasoning using LLMs by means of simplified graph structures as enumerated node adjacency lists and connections [2][3][4], graph tasks such as traversal are inherently reduced to trivial search retrieval problems due to these flattened representations. Hence, the ability of LLMs to navigate over more complex and structured representations is still not well explored. We explore this particular question in-depth in this paper, by going one step further than previous works. Specifically, we prompt LLMs to reason over semi-structured graph representations, instead of plainly verbalized unstructured representations like previous works, while still maintaining the language compatibility of LLMs.  [^3] In this paper’s context, semi-structured representations involve breaking down a structured representation into uni-directional text, while still maintaining some form of element relationship. In our case, adjacency matrix representations involve a tabular relation set, which requires tracking both the column and row origin of a node. This structure introduces complexity to the reasoning task. Particularly, we are interested in exploring the ability in LLMs to handle such representational complexity.